The force applied to the cart by spring scale Fa i


Friction Force is 4.846
Answer #1
as cart moves right friction force acts left
so Fb+ frictional foce = Fa
so Fb = 10.5 - 4.846 = 5.654 N
so reading on Fb = 5.654 N
Answer #2
Using newton's first law of mostion,Net force on a abody is zero if it continues to move with constant velocity.
hence FB = Fa-Friction force(which acts backwards)
= 10.5-4.846
= 5.654 N